### v1

@@
Scientific discovery constitutes an ongoing development of explanatory theories that advance understanding and prediction of reality1, 2; contributing to technological and social progress. Throughout history, our capacity to formulate and test theories about the universe has evolved from early philosophical inquiry (Plato3 and Aristotle4) into modern scientific investigation. Fundamentally, conjecture and criticism form the basis for scientific advancement, where theories face rigorous testing against empirical evidence and competing explanations5, 6. Scientists have achieved remarkable breakthroughs through this approach; however, the scientific endeavour operates within inherent human cognitive constraints. Developing better explanatory theories1, 2 requires researchers to synthesise existing knowledge, formulate testable hypotheses, design rigorous experiments, and interpret results within broader theoretical frameworks. As research output and complexity grow exponentially7, 8 (>>2.8 million/yr since 2022; $\sim$5.6% growth9), researchers face increasing cognitive and practical limitations in their ability to synthesise existing knowledge and discover novel insights10. Individual researchers typically report reading $\sim$200 – 300 articles per year, yet many fields now produce thousands annually, suggesting that complete coverage and awareness of disciplinary literature may no longer be practically achievable11. Hypothesis generation increasingly occurs within narrowed conceptual spaces constrained by specialised training, while experimental design can suffer from limited cross-disciplinary methodological exposure. Such limitations may affect scientists’ capacity to develop unifying explanatory frameworks that transcend disciplinary boundaries and address fundamental questions – the ”better explanations” that can drive scientific progress1, 12.
@@
Computational approaches offer the potential to augment human research capabilities, while alleviating some of these limitations within the traditional scientific process. Early theoretical work in chaos theory13 and computational principles14 laid essential foundations for algorithmic processing of complex scientific problems. Today, AI applications are transforming scientific discovery across multiple domains from chemistry15, synthetic biology16, 17, materials science18, 19, 20, mathematics21, 22, and algorithm development23; accelerating research productivity and discovery in these areas. For example, AlphaFold has enabled accurate structure prediction for many proteins within hours17, substantially accelerating downstream research24. However, these specialised AI systems operate within narrow scientific domains, requiring significant human expertise and/or explicit programming to solve predefined problems. Modern transformer architectures25, combined with large text corpora and scaling of computational infrastructure26, have enabled the development of large language models (LLMs) which demonstrate broad scope versatility, general reasoning, and fluency in scientific discourse across multiple domains27, 28, 29, 30. Agentic frameworks leverage these capabilities by integrating LLMs within autonomous architectures capable of goal-directed planning, tool use, and environmental feedback31. These systems differ fundamentally from previous AI approaches through their capacity to navigate key phases of scientific inquiry within a unified architecture, including hypothesis generation32, 33, 34, experimental design and execution35, 36, 37, manuscript writing38, and paper evaluation39.
@@
The first complete agentic pipeline for in silico computational research38 autonomously generated machine learning research from concept to manuscript, with one such article passing peer review for inclusion at a scientific workshop40. Current end-to-end agentic scientific discovery frameworks continue to operate primarily in simulated or abstract digital environments40, 41, 42, 43, 44, 45, typically producing outcomes with predictive power but lacking explanation of underlying causal mechanisms. While these predictions can inform subsequent theories, the absence of direct mechanistic insight constrains hypothesis development and generalisable principles1, 46. Systems that both generate predictions and transparently articulate causal reasoning would significantly advance autonomous scientific discovery by contributing to explanatory frameworks. Embodied AI offers a pathway to address the limitations of previous (purely computational) approaches, as it integrates perception, action, and reasoning into agents within environments that provide direct feedback47, 48, 49. As a result, the system can test hypotheses via physical experimentation (utilising the foundation provided by online platforms that interact with humans, through to manipulation of robotic appendages35) and collect results1, 46 to support or refute proposed explanatory models. Systems have previously developed such potential in a chemical discovery context by automating certain chemical experimentation tasks (e.g., physically mixing reagents and observing results)50, 51, but these systems lacked the capacity for autonomous hypothesis formulation or refinement of understanding based on experimental results.
@@
Here we explore this potential with the goal of developing an autonomous end-to-end scientific discovery pipeline capable of conducting real-world experiments. To this end, we implemented an agentic system incorporating hypothesis generation through experimental design, physical experimental implementation, data analysis, result interpretation, theory refinement, visualisation, and reporting. Completion of a full scientific study required on average 17 hours’ processing time and averaged a total marginal cost of $\sim$\$114 USD per research project (not including the human participant payments of $\sim$\$4,500 USD for the current experiment). We selected cognitive science as the validation domain based on our expertise and the field’s established frameworks for remote experimentation. Given pre-validated cognitive paradigms testing visual working memory (VWM)52, mental rotation53, and imagery vividness54; we tasked the system with completing three lines of research inquiry. Study 1 conducted a controlled intervention, collecting new data; Studies 2–3 analysed the same 288‑participant dataset from Study 1 with distinct hypotheses and analyses. Study 1 examined whether VWM precision and mental rotation performance share representational constraints, finding no correlation between individual performance patterns despite both tasks showing expected difficulty effects; attributed to the established ”reliability paradox”55. Study 2 explored whether imagery vividness influences how previous stimuli bias current perception and memory, finding that individuals with stronger imagery showed no greater carryover effects between trials, challenging theories that imagery and perception rely on common processing mechanisms. Study 3 investigated whether the precision of VWM predicts broader spatial reasoning abilities, finding negligible relationships and suggesting that apparent connections between visual-spatial tasks reflect general cognitive factors rather than specific shared processes. Together, these results demonstrate the feasibility of AI-driven empirical research, and steps toward an embodied AI framework that navigates all the key components of experimental scientific workflows.
@@ 1 Methods
The end-to-end pipeline includes: (1) a hypothesis formulation engine that identifies potential research questions and testable predictions by searching and validating novelty, breakthrough potential, and feasibility; (2) an experimental protocol engine that designs methodologies, presented as a pre-registration report following Open Science Framework guidelines56, and includes preliminary power analyses as required; (3) an implementation engine, currently interfaced with platforms for cognitive science; (4) a data analysis engine that designs and executes a transparent processing pipeline, covering raw data cleaning, outlier analysis, statistical testing, and interpretation of outcomes; (5) scientific decision-making, specifically synthesising and analysing experimental outcomes through inference frameworks to determine follow-up experiments and/or studies; (6) a visualisation engine which designs and constructs a set of figures and tables to illustrate results collated across experiments; (7) drafting of a complete manuscript incorporating visualisations and validated citations; (8) ‘peer’-style evaluation; and (9) construction of a final formatted manuscript. It achieves a key goal by bridging discovery from in silico computational domains, to the real world, enabling the system to conduct empirical testing of hypotheses with experimental interventions on human participants, and to perform detailed analyses of complex, noisy real-world data. While demonstrated here through cognitive psychology experiments, the architecture employs domain-general principles designed to be applicable across diverse scientific fields and achieves a fundamental goal of the emerging ‘self‑driving‑laboratory’ paradigm57, 58.
@@
The system leverages a hierarchical multi-agent architecture59 to autonomously produce scientific research through the coordinated and collaborative efforts of a task force of specialised AI agents (Figure 1). Unlike deterministic pipelines, each agent functions as an autonomous entity capable of receiving inputs, applying domain-specific reasoning, and producing outputs that advance the investigation. Within the dynamically layered network of agent interactions, consisting of ‘orchestrators’ (agents with the ability to coordinate and create further sub-agents for themselves) and ‘specialists’ (agents excelling in honed skillsets, such as coding, troubleshooting, or review), a single top-level orchestrator (the master agent) coordinates the entire scientific workflow from beginning to end. This structure results in an emergent cascade of expertise and directed attentional flows, key to successfully navigating the unpredictable course of real-world experimentation59, 60.
@@
The framework accommodates multiple operation modes along a continuum from autonomous investigation to human-in-the-loop machine learning collaboration models61. When running autonomously, the system progresses independently from initial hypothesis to completed manuscript, with each agent making decisions guided by disciplinary standards and outputs from preceding stages. Collaborative modes, on the other hand, allow researchers to provide strategic input at specific intervention points, such as suggesting publications for consideration, bounding decision spaces according to their domain expertise or research priorities, specifying methodological constraints, or outlining visualisation preferences. Researchers can also review and refine all outputs produced by the agents at any stage, providing a means of external quality control throughout the process. The system facilitates these collaborative capabilities without sacrificing its end-to-end functionality, ensuring adaptability across varied research contexts and collaboration models. This flexible architecture enables researchers to deploy the system according to their specific needs – from delegating discrete tasks to commissioning complete investigations with minimal supervision.
@@
LLMs exhibit broad capabilities62 yet typically struggle with planning over extended durations and self-verification63, 64. To address these limitations, we established a foundational cognitive control framework for the system comprising four operators derived from psychological science – abstraction, metacognition, decomposition, and autonomy (Figure 2). These operators coordinate planning, tool use, monitoring, evaluation, and refinement across research workflows, serving as computational analogues of the human executive functions that facilitate complex multi-stage inquiries. Each operator draws upon and extends established techniques, combined within the multi-agent system to support empirical investigation with minimal human oversight.
@@
Abstraction. The process of focusing on general patterns rather than instance-specific details65, 66 was operationalised as knowledge induction by enabling agents to develop their own heuristics and instructions rather than constraining them with predetermined directives. Concretely, this involved initial elicitation of latent background premises67, self-driven exploration of problem scope (conceptually related to the ‘Self-Ask’ method68), and automated instruction generation (as validated previously69). By beginning with universal principles, the system maintains a broader conceptual search space for potential scientific insights. This implementation mirrors how human scientists maintain conceptual flexibility when developing novel theories70, allowing exploration across disciplinary boundaries that might otherwise be constrained by specialised training.
@@
Metacognition. Awareness and regulation of one’s own thinking processes71, 72 was operationalised at two levels, individual and collective, to assess and refine agent reasoning. While frontier LLMs inherently employ forms of internal test-time compute that dynamically scale with task complexity, the system implements explicit self-evaluation protocols that assess evidence quality, logical coherence, and rigour. At the individual level, this was implemented through self-reflective chains of thought73, enabling each agent to interrogate its underlying assumptions prior to reaching conclusions. At the collective level, agent groups developed awareness of their joint thinking through a reflective process operating on all agents’ reasoning traces, similar to ‘Tree of Thoughts’ inference74, but across several different agents and utilising an external ‘Agent-as-a-Judge’75 to assess, refine, and arbitrate those traces to align the group on a decided path. These structured self-reflection mechanisms enhance accuracy in complex reasoning tasks73, 76 and facilitate transparent documentation of the evaluation processes.
@@
Decomposition. The breaking down of complex problems into more manageable components77, 78 – was operationalised in the framework as explicit structuring of the solution search space. This decomposition enhances the system’s capacity to manage the intricacy of multi-stage scientific workflows while maintaining precision at each step. Specifically, parameterisation of logical reasoning steps, conceptually aligned with ‘least-to-most’ prompting79, identifies constituent task components. This improves the tractability and transparency of multi-stage scientific workflows and enables verification and refinement of each component to maintain step-level precision. In addition, the system’s recursive divide-and-conquer agentic architecture facilitates on-demand subdivision of effort as required, providing flexibility to adapt to challenging tasks80, thereby improving reliability in the production of the required scientific deliverables.
@@
Autonomy. Self-directed goal pursuit81, 82 was implemented in orchestrator agents as local decision-making on tasks, constrained by explicit system objectives. Each orchestrator independently works to complete assigned goals by iterating through a propose-validate-refine process akin to the ‘Self-Refine’ algorithm83. Iteration was governed by three policies: initiation, replanning, and termination. Initiation of sub-agents was invoked upon request of the agent to assist with task completion. Replanning was triggered when validation failed or marginal improvement on acceptance tests fell below a patience threshold. Termination occurred either when validation and quality acceptance checks were all satisfied, or when pre-defined recursion limits were exhausted. Within each iteration, validation of the agents’ proposition utilised appropriate external tools where possible to assist with self-correction (shown previously to improve performance84). These task feedback signals were then synthesised using a framework analogous to ‘Reflexion’85 and incorporated by the agent to inform subsequent propositions. The iterative self-editing continued until the stopping rules were met, at which point the agent returned its work to its orchestrator.
@@
Humans navigate complex tasks utilising sophisticated memory systems that can: i) hold and manipulate information in working memory86, 87, ii) selectively filter relevant details for the task at hand88, 89, and iii) offload information to external resources when internal capacity is exceeded90. For the multi-agent system to maintain coherence over long periods, we emulated these capabilities in the system through the implementation of a dynamic Retrieval-Augmented Generation (d-RAG) system (comparable to ‘DRAGIN’91) and construction of retrievable artifacts. Extending existing RAG architectures92, the d-RAG provides dynamic memory which augments each agent with the cognitive flexibility, prior knowledge, and specific information necessary to carry out its task. Instead of using the same reference material regardless of context, the d-RAG creates and evolves specialised knowledge repositories for each research direction traversed. It functions analogously to how researchers develop domain-specific expertise through targeted literature engagement, and access knowledge during scientific inquiry, combining working memory with cognitive offloading to external resources. The d-RAG forms the core of a multi-tier ’search engine’ tool accessible to the specialist archivist agent, which assists orchestrators requiring real-world information, reducing reliance solely on trained LLM knowledge that may be prone to factual inconsistencies93. The search engine facilitates three depths of inquiry: (1) broad academic database searches via the APIs for Semantic Scholar94, OpenAlex95, and PubMed, (2) within-text multi-article d-RAG queries (akin to ‘PaperQA-2’96), and (3) paper-specific question-answering. The system progressively builds its knowledge base by processing retrieved academic papers in response to agents’ queries, discarding irrelevant retrievals to maintain focused knowledge representations tailored to the specific research question. In addition, by allowing the system to offload complexity to specialised components and file artifacts, a concise and compact representation of the overall research state can be maintained at all points.
@@
To increase robustness of the system across the various scientific tasks and unique challenges posed by each, we employed a Mixture of Agents (MoA) approach leveraging complementary strengths of different frontier LLMs97. Though MoA may become less necessary as model intelligence advances, current training data choices and reinforcement learning by human feedback processes predispose models to inherent biases that can limit a model’s scientific utility. As such, the best performing models for specific aspects of the system were selected to operate conjointly, enabling them to collaboratively accomplish complex tasks that proved highly challenging for any individual model to complete alone. The frontier models utilised include Anthropic’s Claude 4 Sonnet, OpenAI’s o3-mini & o1, xAI’s Grok-3, Mistral’s Pixtral Large, and Google’s Gemini 2.5 Pro.
@@
Frontier LLMs encode high-dimensional representations of knowledge through training on vast text corpora98, enabling interpolation across distant scientific domains62 and disparate concepts – such as quantum entanglement and photosynthetic energy transfer – to produce inquiries that may lie in unexplored interstices of existing research. Recent evidence of emergent reasoning in LLMs99 and their ability to blend conceptual domains100 supports this notion, suggesting that latent spaces encode abstractions conducive to creative synthesis. The vector-space abstraction enabling transfer in LLMs may be functionally analogous – though not necessarily architecturally identical – to the neural abstraction supporting human task generalisation101. Seminal work in psychology suggests transfer between learning contexts requires shared structural elements102, 103 and exemplified the notion of abstraction in cognition104, 105, demonstrating how mental schemas facilitate knowledge transfer across disparate domains. Similarly, LLMs may interpolate between abstracted representations of their training data to produce novel information in new contexts other than the original domains106, 107. Although LLM-generated ideas are novel in that they are yet to be explored in human documentation, they fundamentally represent an amalgam of current human knowledge and are thus constrained by the models’ training data108. The system here overcomes this limitation through empirical hypothesis testing, as each completed research idea serves as a basis from which the system can ideate further scientific enquiry (Figure 3). By establishing a recursive learning lineage that continually extends its knowledge beyond the original LLM training corpus, it facilitates the potential to yield truly novel hypotheses.

### v2

@@
Scientific discovery constitutes an ongoing development of explanatory theories that advance understanding and prediction of realitydeutsch_beginning_2011, popper_conjectures_1963; contributing to technological and social progress. Throughout history, our capacity to formulate and test theories about the universe has evolved from early philosophical inquiry (Platoplato_republic_2013 and Aristotlearistotle_posterior_1960) into modern scientific investigation. Fundamentally, conjecture and criticism form the basis for scientific advancement, where theories face rigorous testing against empirical evidence and competing explanationsdeutsch_fabric_1997, popper_logic_1959. Scientists have achieved remarkable breakthroughs through this approach; however, the scientific endeavour operates within inherent human cognitive constraints. Developing better explanatory theoriesdeutsch_beginning_2011, popper_conjectures_1963 requires researchers to synthesise existing knowledge, formulate testable hypotheses, design rigorous experiments, and interpret results within broader theoretical frameworks. As research output and complexity grow exponentiallybornmann_growth_2015, landhuis_scientific_2016 (>>2.8 million/yr since 2022; $\sim$5.6% growthhanson_strain_2024), researchers face increasing cognitive and practical limitations in their ability to synthesise existing knowledge and discover novel insightstenopir_scholarly_2015. Individual researchers typically report reading $\sim$200 – 300 articles per year, yet many fields now produce thousands annually, suggesting that complete coverage and awareness of disciplinary literature may no longer be practically achievabletenopir_scholarly_2012. Hypothesis generation increasingly occurs within narrowed conceptual spaces constrained by specialised training, while experimental design can suffer from limited cross-disciplinary methodological exposure. Such limitations may affect scientists’ capacity to develop unifying explanatory frameworks that transcend disciplinary boundaries and address fundamental questions – the ”better explanations” that can drive scientific progressdeutsch_beginning_2011, kuhn_structure_1962.
@@
Computational approaches offer the potential to augment human research capabilities, while alleviating some of these limitations within the traditional scientific process. Early theoretical work in chaos theorylorenz_deterministic_1963 and computational principlesshannon_probabilistic_1956 laid essential foundations for algorithmic processing of complex scientific problems. Today, AI applications are transforming scientific discovery across multiple domains from chemistrybuchanan_dendral_1981, synthetic biologyhayes_simulating_2025, jumper_highly_2021, materials sciencemerchant_scaling_2023, pyzer-knapp_accelerating_2022, szymanski_autonomous_2023, mathematicslenat_automated_1977, romera-paredes_mathematical_2024, and algorithm developmentfawzi_discovering_2022; accelerating research productivity and discovery in these areas. For example, AlphaFold has enabled accurate structure prediction for many proteins within hoursjumper_highly_2021, substantially accelerating downstream researchtunyasuvunakool_highly_2021. However, these specialised AI systems operate within narrow scientific domains, requiring significant human expertise and/or explicit programming to solve predefined problems. Modern transformer architecturesvaswani_attention_2017, combined with large text corpora and scaling of computational infrastructureopenai_gpt-4_2023, have enabled the development of large language models (LLMs) which demonstrate broad scope versatility, general reasoning, and fluency in scientific discourse across multiple domainsopenai_openai_2025, deepseek-ai_deepseek-r1_2025, anthropic_system_2025, google_gemini_2025. Agentic frameworks leverage these capabilities by integrating LLMs within autonomous architectures capable of goal-directed planning, tool use, and environmental feedbackwang_survey_2024. These systems differ fundamentally from previous AI approaches through their capacity to navigate key phases of scientific inquiry within a unified architecture, including hypothesis generationborrego_research_2025, ghafarollahi_sciagents_2025, gottweis_towards_2025, experimental design and executionburger_mobile_2020, ghafarollahi_atomagents_2024, m_bran_augmenting_2024, manuscript writinglu_ai_2024, and paper evaluationhuang_papereval_2025.
@@
The first complete agentic pipeline for in silico computational researchlu_ai_2024 autonomously generated machine learning research from concept to manuscript, with one such article passing peer review for inclusion at a scientific workshopyamada_ai_2025. Current end-to-end agentic scientific discovery frameworks continue to operate primarily in simulated or abstract digital environmentsyamada_ai_2025, intology_zochi_2025, schmidgall_agent_2025, swanson_virtual_2025, weng_cycleresearcher_2024, ifargan_autonomous_2025, typically producing outcomes with predictive power but lacking explanation of underlying causal mechanisms. While these predictions can inform subsequent theories, the absence of direct mechanistic insight constrains hypothesis development and generalisable principlesdeutsch_beginning_2011, pearl_causality_2009. Systems that both generate predictions and transparently articulate causal reasoning would significantly advance autonomous scientific discovery by contributing to explanatory frameworks. Embodied AI offers a pathway to address the limitations of previous (purely computational) approaches, as it integrates perception, action, and reasoning into agents within environments that provide direct feedbackbrooks_intelligence_1991, clark_being_1996, pfeifer_how_2007. As a result, the system can test hypotheses via physical experimentation (utilising the foundation provided by online platforms that interact with humans, through to manipulation of robotic appendagesburger_mobile_2020) and collect resultsdeutsch_beginning_2011, pearl_causality_2009 to support or refute proposed explanatory models. Systems have previously developed such potential in a chemical discovery context by automating certain chemical experimentation tasks (e.g., physically mixing reagents and observing results)pagel_validation_2024, boiko_autonomous_2023, but these systems lacked the capacity for autonomous hypothesis formulation or refinement of understanding based on experimental results.
@@
Here we explore this potential with the goal of developing an autonomous end-to-end scientific discovery pipeline capable of conducting real-world experiments. To this end, we implemented an agentic system incorporating hypothesis generation through experimental design, physical experimental implementation, data analysis, result interpretation, theory refinement, visualisation, and reporting. Completion of a full scientific study required on average 17 hours’ processing time and averaged a total marginal cost of $\sim$\$114 USD per research project (not including the human participant payments of $\sim$\$4,500 USD for the current experiment). We selected cognitive science as the validation domain based on our expertise and the field’s established frameworks for remote experimentation. Given pre-validated cognitive paradigms testing visual working memory (VWM)zhang_discrete_2008, mental rotationshepard_mental_1971, and imagery vividnessmarks_new_1995; we tasked the system with completing three lines of research inquiry. Study 1 conducted a controlled intervention, collecting new data; Studies 2–3 analysed the same 288‑participant dataset from Study 1 with distinct hypotheses and analyses. Study 1 examined whether VWM precision and mental rotation performance share representational constraints, finding no correlation between individual performance patterns despite both tasks showing expected difficulty effects; attributed to the established ”reliability paradox”hedge_reliability_2018. Study 2 explored whether imagery vividness influences how previous stimuli bias current perception and memory, finding that individuals with stronger imagery showed no greater carryover effects between trials, challenging theories that imagery and perception rely on common processing mechanisms. Study 3 investigated whether the precision of VWM predicts broader spatial reasoning abilities, finding negligible relationships and suggesting that apparent connections between visual-spatial tasks reflect general cognitive factors rather than specific shared processes. Together, these results demonstrate the feasibility of AI-driven empirical research, and steps toward an embodied AI framework that navigates all the key components of experimental scientific workflows.
@@ 1 Methods
The end-to-end pipeline includes: (1) a hypothesis formulation engine that identifies potential research questions and testable predictions by searching and validating novelty, breakthrough potential, and feasibility; (2) an experimental protocol engine that designs methodologies, presented as a pre-registration report following Open Science Framework guidelinesfoster_open_2017, and includes preliminary power analyses as required; (3) an implementation engine, currently interfaced with platforms for cognitive science; (4) a data analysis engine that designs and executes a transparent processing pipeline, covering raw data cleaning, outlier analysis, statistical testing, and interpretation of outcomes; (5) scientific decision-making, specifically synthesising and analysing experimental outcomes through inference frameworks to determine follow-up experiments and/or studies; (6) a visualisation engine which designs and constructs a set of figures and tables to illustrate results collated across experiments; (7) drafting of a complete manuscript incorporating visualisations and validated citations; (8) ‘peer’-style evaluation; and (9) construction of a final formatted manuscript. It achieves a key goal by bridging discovery from in silico computational domains, to the real world, enabling the system to conduct empirical testing of hypotheses with experimental interventions on human participants, and to perform detailed analyses of complex, noisy real-world data. While demonstrated here through cognitive psychology experiments, the architecture employs domain-general principles designed to be applicable across diverse scientific fields and achieves a fundamental goal of the emerging ‘self‑driving‑laboratory’ paradigmbayley_autonomous_2024, lo_review_2024.
@@
The system leverages a hierarchical multi-agent architecturezhang_agentorchestra_2025 to autonomously produce scientific research through the coordinated and collaborative efforts of a task force of specialised AI agents (Figure 1). Unlike deterministic pipelines, each agent functions as an autonomous entity capable of receiving inputs, applying domain-specific reasoning, and producing outputs that advance the investigation. Within the dynamically layered network of agent interactions, consisting of ‘orchestrators’ (agents with the ability to coordinate and create further sub-agents for themselves) and ‘specialists’ (agents excelling in honed skillsets, such as coding, troubleshooting, or review), a single top-level orchestrator (the master agent) coordinates the entire scientific workflow from beginning to end. This structure results in an emergent cascade of expertise and directed attentional flows, key to successfully navigating the unpredictable course of real-world experimentationzhang_agentorchestra_2025, fourney_magentic-one_2024.
@@
The framework accommodates multiple operation modes along a continuum from autonomous investigation to human-in-the-loop machine learning collaboration modelsmosqueira-rey_human---loop_2023. When running autonomously, the system progresses independently from initial hypothesis to completed manuscript, with each agent making decisions guided by disciplinary standards and outputs from preceding stages. Collaborative modes, on the other hand, allow researchers to provide strategic input at specific intervention points, such as suggesting publications for consideration, bounding decision spaces according to their domain expertise or research priorities, specifying methodological constraints, or outlining visualisation preferences. Researchers can also review and refine all outputs produced by the agents at any stage, providing a means of external quality control throughout the process. The system facilitates these collaborative capabilities without sacrificing its end-to-end functionality, ensuring adaptability across varied research contexts and collaboration models. This flexible architecture enables researchers to deploy the system according to their specific needs – from delegating discrete tasks to commissioning complete investigations with minimal supervision.
@@
LLMs exhibit broad capabilitiesbubeck_sparks_2023 yet typically struggle with planning over extended durations and self-verificationstechly_self-verification_2024, valmeekam_llms_2024. To address these limitations, we established a foundational cognitive control framework for the system comprising four operators derived from psychological science – abstraction, metacognition, decomposition, and autonomy (Figure 2). These operators coordinate planning, tool use, monitoring, evaluation, and refinement across research workflows, serving as computational analogues of the human executive functions that facilitate complex multi-stage inquiries. Each operator draws upon and extends established techniques, combined within the multi-agent system to support empirical investigation with minimal human oversight.
@@
Abstraction. The process of focusing on general patterns rather than instance-specific detailsgentner_analogy_2017, goldstone_transfer_2005 was operationalised as knowledge induction by enabling agents to develop their own heuristics and instructions rather than constraining them with predetermined directives. Concretely, this involved initial elicitation of latent background premisesliu_generated_2021, self-driven exploration of problem scope (conceptually related to the ‘Self-Ask’ methodpress_measuring_2022), and automated instruction generation (as validated previouslyzhou_large_2022). By beginning with universal principles, the system maintains a broader conceptual search space for potential scientific insights. This implementation mirrors how human scientists maintain conceptual flexibility when developing novel theoriesdunbar_scientific_2012, allowing exploration across disciplinary boundaries that might otherwise be constrained by specialised training.
@@
Metacognition. Awareness and regulation of one’s own thinking processesfleming_self-evaluation_2017, shea_supra-personal_2014 was operationalised at two levels, individual and collective, to assess and refine agent reasoning. While frontier LLMs inherently employ forms of internal test-time compute that dynamically scale with task complexity, the system implements explicit self-evaluation protocols that assess evidence quality, logical coherence, and rigour. At the individual level, this was implemented through self-reflective chains of thoughtwei_chain--thought_2022, enabling each agent to interrogate its underlying assumptions prior to reaching conclusions. At the collective level, agent groups developed awareness of their joint thinking through a reflective process operating on all agents’ reasoning traces, similar to ‘Tree of Thoughts’ inferenceyao_tree_2023, but across several different agents and utilising an external ‘Agent-as-a-Judge’zhuge_agent-as--judge_2024 to assess, refine, and arbitrate those traces to align the group on a decided path. These structured self-reflection mechanisms enhance accuracy in complex reasoning taskswei_chain--thought_2022, renze_self-reflection_2024 and facilitate transparent documentation of the evaluation processes.
@@
Decomposition. The breaking down of complex problems into more manageable componentsanderson_architecture_1983, newell_human_1972 – was operationalised in the framework as explicit structuring of the solution search space. This decomposition enhances the system’s capacity to manage the intricacy of multi-stage scientific workflows while maintaining precision at each step. Specifically, parameterisation of logical reasoning steps, conceptually aligned with ‘least-to-most’ promptingzhou_least--most_2022, identifies constituent task components. This improves the tractability and transparency of multi-stage scientific workflows and enables verification and refinement of each component to maintain step-level precision. In addition, the system’s recursive divide-and-conquer agentic architecture facilitates on-demand subdivision of effort as required, providing flexibility to adapt to challenging tasksprasad_adapt_2024, thereby improving reliability in the production of the required scientific deliverables.
@@
Autonomy. Self-directed goal pursuitbandura_social_2001, ryan_self-determination_2000 was implemented in orchestrator agents as local decision-making on tasks, constrained by explicit system objectives. Each orchestrator independently works to complete assigned goals by iterating through a propose-validate-refine process akin to the ‘Self-Refine’ algorithmmadaan_self-refine_2023. Iteration was governed by three policies: initiation, replanning, and termination. Initiation of sub-agents was invoked upon request of the agent to assist with task completion. Replanning was triggered when validation failed or marginal improvement on acceptance tests fell below a patience threshold. Termination occurred either when validation and quality acceptance checks were all satisfied, or when pre-defined recursion limits were exhausted. Within each iteration, validation of the agents’ proposition utilised appropriate external tools where possible to assist with self-correction (shown previously to improve performancegou_critic_2023). These task feedback signals were then synthesised using a framework analogous to ‘Reflexion’shinn_reflexion_2023 and incorporated by the agent to inform subsequent propositions. The iterative self-editing continued until the stopping rules were met, at which point the agent returned its work to its orchestrator.
@@
Humans navigate complex tasks utilising sophisticated memory systems that can: i) hold and manipulate information in working memorybaddeley_working_1974, desposito_cognitive_2015, ii) selectively filter relevant details for the task at handdesimone_neural_1995, corbetta_control_2002, and iii) offload information to external resources when internal capacity is exceededrisko_cognitive_2016. For the multi-agent system to maintain coherence over long periods, we emulated these capabilities in the system through the implementation of a dynamic Retrieval-Augmented Generation (d-RAG) system (comparable to ‘DRAGIN’su_dragin_2024) and construction of retrievable artifacts. Extending existing RAG architectureslewis_retrieval-augmented_2020, the d-RAG provides dynamic memory which augments each agent with the cognitive flexibility, prior knowledge, and specific information necessary to carry out its task. Instead of using the same reference material regardless of context, the d-RAG creates and evolves specialised knowledge repositories for each research direction traversed. It functions analogously to how researchers develop domain-specific expertise through targeted literature engagement, and access knowledge during scientific inquiry, combining working memory with cognitive offloading to external resources. The d-RAG forms the core of a multi-tier ’search engine’ tool accessible to the specialist archivist agent, which assists orchestrators requiring real-world information, reducing reliance solely on trained LLM knowledge that may be prone to factual inconsistencieshuang_survey_2025. The search engine facilitates three depths of inquiry: (1) broad academic database searches via the APIs for Semantic Scholarkinney_semantic_2023, OpenAlexpriem_openalex_2022, and PubMed, (2) within-text multi-article d-RAG queries (akin to ‘PaperQA-2’skarlinski_language_2024), and (3) paper-specific question-answering. The system progressively builds its knowledge base by processing retrieved academic papers in response to agents’ queries, discarding irrelevant retrievals to maintain focused knowledge representations tailored to the specific research question. In addition, by allowing the system to offload complexity to specialised components and file artifacts, a concise and compact representation of the overall research state can be maintained at all points.
@@
To increase robustness of the system across the various scientific tasks and unique challenges posed by each, we employed a Mixture of Agents (MoA) approach leveraging complementary strengths of different frontier LLMswang_mixture--agents_2024. Though MoA may become less necessary as model intelligence advances, current training data choices and reinforcement learning by human feedback processes predispose models to inherent biases that can limit a model’s scientific utility. As such, the best performing models for specific aspects of the system were selected to operate conjointly, enabling them to collaboratively accomplish complex tasks that proved highly challenging for any individual model to complete alone. The frontier models utilised include Anthropic’s Claude 4 Sonnet, OpenAI’s o3-mini & o1, xAI’s Grok-3, Mistral’s Pixtral Large, and Google’s Gemini 2.5 Pro.
@@
Frontier LLMs encode high-dimensional representations of knowledge through training on vast text corporabommasani_opportunities_2022, enabling interpolation across distant scientific domainsbubeck_sparks_2023 and disparate concepts – such as quantum entanglement and photosynthetic energy transfer – to produce inquiries that may lie in unexplored interstices of existing research. Recent evidence of emergent reasoning in LLMswei_emergent_2022 and their ability to blend conceptual domainsfauconnier_way_2002 supports this notion, suggesting that latent spaces encode abstractions conducive to creative synthesis. The vector-space abstraction enabling transfer in LLMs may be functionally analogous – though not necessarily architecturally identical – to the neural abstraction supporting human task generalisationgarner_knowledge_2023. Seminal work in psychology suggests transfer between learning contexts requires shared structural elementsthorndike_educational_1903, woodworth_influence_1901 and exemplified the notion of abstraction in cognitionbartlett_remembering_1932, piaget_origins_1952, demonstrating how mental schemas facilitate knowledge transfer across disparate domains. Similarly, LLMs may interpolate between abstracted representations of their training data to produce novel information in new contexts other than the original domainskojima_large_2022, zhou_hypothesis_2024. Although LLM-generated ideas are novel in that they are yet to be explored in human documentation, they fundamentally represent an amalgam of current human knowledge and are thus constrained by the models’ training datafelin_theory_2024. The system here overcomes this limitation through empirical hypothesis testing, as each completed research idea serves as a basis from which the system can ideate further scientific enquiry (Figure 3). By establishing a recursive learning lineage that continually extends its knowledge beyond the original LLM training corpus, it facilitates the potential to yield truly novel hypotheses.